# WebArbiter: A Principle-Guided Reasoning Process Reward Model for Web Agents

### v1

@@ Abstract
Web agents hold great potential for automating complex computer tasks, yet their interactions involve long-horizon, sequential decision-making with irreversible actions. In such settings, outcome-based supervision is sparse and delayed, often rewarding incorrect trajectories and failing to support inference-time scaling. This motivates the use of Process Reward Models (WebPRMs) for web navigation, but existing approaches remain limited: scalar WebPRMs collapse progress into coarse, weakly grounded signals, while checklist-based WebPRMs rely on brittle template matching that fails under layout or semantic changes and often mislabels superficially correct actions as successful, providing little insight or interpretability. To address these challenges, we introduce WebArbiter, a reasoning-first, principle-inducing WebPRM that formulates reward modeling as text generation, producing structured justifications that conclude with a preference verdict and identify the action most conducive to task completion under the current context. Training follows a two-stage pipeline: reasoning distillation equips the model with coherent principle-guided reasoning, and reinforcement learning corrects teacher biases by directly aligning verdicts with correctness, enabling stronger generalization. To support systematic evaluation, we release WebPRMBench, a comprehensive benchmark spanning four diverse web environments with rich tasks and high-quality preference annotations. On WebPRMBench, WebArbiter-7B outperforms the strongest baseline, GPT-5, by 9.1 points. In reward-guided trajectory search on WebArena-Lite, it surpasses the best prior WebPRM by up to 7.2 points, underscoring its robustness and practical value in real-world complex web tasks.
@@ 1 Introduction
As shown in Fig. 1, experiments on WebPRMBench show that WebArbiter achieves the highest Avg. BoN Acc among all evaluated models, outperforming the strongest proprietary LLM baseline, GPT-5, by 9.1 points, and consistently surpassing the previous SOTA WebPRM, WebShepherd, across all environments. Beyond static evaluation, WebArbiter also proves effective in practice: in reward-guided trajectory search on WebArena-Lite (Liu et al., 2024b), it delivers substantial gains, surpassing WebShepherd by up to 7.2 points, further demonstrating robustness in realistic interaction settings.
@@
    We show that WebArbiter achieves SOTA performance on WebPRMBench, surpassing both proprietary LLMs and the previous SOTA WebPRM. WebArbiter delivers up to 7.2% gains in reward-guided trajectory search on WebArena-Lite.
@@ 3 Methodology
In this section, we present the design of WebArbiter. Fig. 2 provides an overview of the WebArbiter framework, including the two-stage training pipeline and the inference-time principle-guided decision process. We begin by framing web navigation as a Partially Observable Markov Decision Process (POMDP) in §3.1, then describe how we construct a pairwise-preference dataset for training in §3.2. We introduce the training pipeline of WebArbiter model in §3.3. For clarity, we summarize all notations in Appendix A.
@@ 3.1 Background
We formalize web navigation as a POMDP. The environment $\mathcal{E}$ is defined by a state space $\mathcal{S}$, an action space $\mathcal{A}$, and an observation space $\mathcal{O}$. $T:\mathcal{S}\!\times\!\mathcal{A}\!\to\!\mathcal{S}$ denotes the state transition function. At step $p$, the agent receives a partial observation $o_{p}\!\in\!\mathcal{O}$, executes $a_{p}\!\in\!\mathcal{A}$, and transitions to $s_{p+1}=T(s_{p},a_{p})$ with a new observation $o_{p+1}$. Following WebArena (Zhou et al., 2024), we represent observations using accessibility trees, i.e., text-only encodings of visible interactive elements and their attributes. Given a task instruction $\mathcal{I}$ and the initial state $s_{0}\in\mathcal{S}$, the agent aims to generate a trajectory $\tau=(a_{1},\dots,a_{P})$ that completes the task. Here $P$ is the trajectory length and $a_{p}\in\mathcal{A}$ denotes the action at step $p$. The task evaluator determines whether the task is completed based on the final state.
@@ 4 WebPRMBench
Models Mind2Web WebArena AssistantBench WorkArena Total
@@ 4.1 Benchmark Construction
For each state, the action from the successful trajectory is retained as the positive label, and four rejected alternatives with associated reasoning traces are synthesized to form preference pairs. To ensure data quality, we sample negatives from diverse policy models to broaden coverage, apply rule-based filters to remove invalid or mismatched actions, discard inconsistent cases, and conduct expert verification to further ensure reliability. We also conduct targeted auditing to eliminate potential false negatives. To avoid positional bias, the positive action is not fixed to a specific side and may appear on either side of the preference pair. Reasoning traces are truncated to a fixed length to minimize formatting noise. The resulting benchmark spans 1,150 step-level preference instances across four environments, as shown in Tab. 1. Full construction details and benchmark statistics are provided in Appendix E.
@@ 5.1.2 Main Results
WebArbiter Significantly Outperforms Baselines. As shown in Tab. 2, WebArbiter achieves the highest Avg. Pairwise Acc and Avg. BoN Acc, surpassing both proprietary and open-source LLMs. While LLM-as-judge methods often maintain moderate Pairwise Acc, their performance drops sharply on BoN Acc, revealing poor robustness to hard negatives. In contrast, WebArbiter sustains strong results on both metrics, establishing its reliability under realistic multi-candidate settings. We further analyze inference-time scaling behavior in Appendix G.
Advantage over the SOTA WebPRM. WebShepherd (Chae et al., 2025) represents the previous SOTA WebPRMs. Trained on the same WEBPRM Collection, which was drawn from the Mind2Web environment, WebArbiter-7B achieves an Avg. BoN Acc of 74.60%, surpassing WebShepherd-8B by an absolute gain of 31%. Unlike WebShepherd, which relies on fragile checklists, WebArbiter employs principle-guided reasoning, yielding judgments robust to environment and page variations. Case studies illustrating these differences are provided in Appendix H.
@@ 5.2.1 Experimental Setup and Implementations
Reward-guided trajectory search represents one of the most practical applications of PRMs, as it directly leverages fine-grained step-level supervision to improve decision quality during agent execution. To evaluate WebArbiter in this setting, we conduct experiments on WebArena-Lite [^1] We did not have access to the MAP domain during this work and therefore excluded it from our experiments. (Liu et al., 2024b), which contains diverse, long-horizon tasks such as online shopping and content management, closely reflecting real-world web activities. Performance is measured with Success Rate. Following WebShepherd (Chae et al., 2025), we adopt a Best-of-N sampling strategy: the policy model generates $N=5$ candidate actions for each step, and WebArbiter selects the most promising one through a Knockout Tournament mechanism (Guo et al., 2025b). We evaluate two policies, GPT-4o-mini (OpenAI, 2024a) and GPT-4o (OpenAI, 2024b).
@@ 5.2.2 Downstream Analysis across Domains
As shown in Tab. 5, WebArbiter achieves substantial average improvements under both policy models, significantly outperforming all baselines. These gains stem from two main factors. First, reasoning mitigates spurious correlations that often mislead WebPRMs in domains such as Shopping and Reddit. The improvements on Shopping are particularly pronounced, as these tasks require dense semantic retrieval and inference: stronger policies can propose more promising candidate actions, and WebArbiter’s structured reward modeling further amplifies these advantages. Second, in GitLab, tasks frequently admit multiple equivalent paths. WebShepherd is brittle under such variability, whereas WebArbiter reasons over historical trajectories and the current state to evaluate action validity, enabling stronger generalization in dynamic workflows. We provide qualitative case studies in Appendix H to further illustrate these failure modes of checklist-based supervision. By contrast, CMS exhibits a more template-driven structure, where actions closely follow standardized patterns. In such settings, checklist-based supervision remains comparatively effective, which narrows the relative performance gap. Overall, WebArbiter’s reasoning-first design consistently provides robust, interpretable, and scalable supervision across diverse domains.
@@
Policy WebPRM Shopping CMS Reddit GitLab Avg. $\Delta$
GPT-4o-mini w/o Trajectory Search* 21.74 22.86 19.05 34.38 24.51
GPT-4o-mini 24.44 22.86 26.32 33.33 26.74 +2.23
WebShepherd-8B* 26.09 45.71 23.81 40.62 34.06 +9.55
$\bigstar$ WebArbiter-7B 37.78 42.86 36.84 46.67 41.04 +19.13
GPT-4o w/o Trajectory Search* 23.91 31.43 28.57 56.25 35.04
GPT-4o-mini 26.67 37.14 42.11 40.00 36.48 +1.44
WebShepherd-8B* 30.43 42.86 47.62 46.88 41.95 +6.91
$\bigstar$ WebArbiter-7B 44.44 42.86 52.63 56.67 49.15 +14.11
@@ 6 Conclusion
We presented WebArbiter, a reasoning-first, principle-inducing process reward model that frames reward modeling as structured text generation and produces auditable step-level judgments with rationales. Through reasoning distillation and RL, WebArbiter converts superficial correlations into robust, progress-aware signals that verify genuine task advancement, yield consistent step-level judgments across trajectories, and generalize across dynamic web environments. To support systematic evaluation, we released WebPRMBench, the first comprehensive evaluation benchmark spanning diverse environments for WebPRMs in web navigation, covering four domains with diverse tasks and step-level preference annotations. Extensive experiments demonstrate SOTA performance on WebPRMBench and substantial improvements in reward-guided trajectory search on WebArena-Lite, establishing principle-guided reasoning WebPRMs as a robust and interpretable foundation for scalable web agents.
@@ Appendix A Notation Summary
  • $s_{p},o_{p},a_{p}$: state, observation, and action at step $p$.
@@
  • $\mathcal{D}_{\text{Bench}}$: evaluation dataset for WebArbiter-Bench.
@@
  • $Q$: number of distractors in WebArbiter-Bench (set to $4$).
@@ Appendix C Training Details
All training is conducted on 8 NVIDIA A100-80GB GPUs with fixed random seeds. Our training framework is bead on LLama-Factory (Zheng et al., 2024) and VERL (Sheng et al., 2024)
Distillation Stage. We train the model for 5 epochs with a learning rate of 8e-4, using LoRA with a rank of 128. We apply a cosine learning rate scheduler with a warmup ratio of 0.1. We set the batch size to 256 and the maximum sequence length to 8,192 tokens.
RLVR Stage. We employ the VERL framework for GRPO training. The learning rate is set to 7e-6 for the 7B model, and 9e-6 for the 3B variant. The training uses a fixed batch size of 512 with a mini-batch size of 128, and adopts Fully Sharded Data Parallel (FSDP) for enhanced memory efficiency. For rollout generation, we deploy vLLM with tensor parallelism of 4 and GPU memory utilization limited to 0.4. Response sampling uses standard parameters (temperature=1.0, top-p=1.0), generating 7 candidate responses per prompt. We apply KL regularization with a coefficient of $1.0\times 10^{-3}$ and a clip ratio of 0.2. The maximum input sequence length is 8,192 tokens, and the maximum response length is 4,096 tokens.
@@ E.2 Dataset Composition and Statistics
Appendix G Inference-Time Scaling

### v2

@@ Abstract
Web agents hold great potential for automating complex computer tasks, yet their interactions involve long-horizon, sequential decision-making with irreversible actions. In such settings, outcome-based supervision is sparse and delayed, often rewarding incorrect trajectories and failing to support inference-time scaling. This motivates the use of Process Reward Models (WebPRMs) for web navigation, but existing approaches remain limited: scalar WebPRMs collapse progress into coarse, weakly grounded signals, while checklist-based WebPRMs rely on brittle template matching that fails under layout or semantic changes and often mislabels superficially correct actions as successful, providing little insight or interpretability. To address these challenges, we introduce WebArbiter, a reasoning-first, principle-inducing WebPRM that formulates reward modeling as text generation, producing structured justifications that conclude with a preference verdict and identify the action most conducive to task completion under the current context. Training follows a two-stage pipeline: reasoning distillation equips the model with coherent principle-guided reasoning, and reinforcement learning corrects teacher biases by directly aligning verdicts with correctness, enabling stronger generalization. To support systematic evaluation, we release WebPRMBench, a comprehensive benchmark spanning four diverse web environments with rich tasks and high-quality preference annotations. On WebPRMBench, WebArbiter-7B outperforms the strongest baseline, GPT-5, by 9.1 points. In reward-guided trajectory search on WebArena-Lite, it surpasses the best prior WebPRM by up to 6.4 points, underscoring its robustness and practical value in complex web tasks.
@@ 1 Introduction
As shown in Fig. 1, experiments on WebPRMBench show that WebArbiter achieves the highest Avg. BoN Acc among all evaluated models, outperforming the strongest proprietary LLM baseline, GPT-5, by 9.1 points, and consistently surpassing the previous SOTA WebPRM, WebShepherd, across all environments. Beyond static evaluation, WebArbiter also proves effective in practice: in reward-guided trajectory search on WebArena-Lite (Liu et al., 2024b), it delivers substantial gains, surpassing WebShepherd by up to 6.4 points, further demonstrating robustness in realistic interaction settings.
@@
    We show that WebArbiter achieves SOTA performance on WebPRMBench, surpassing both proprietary LLMs and the previous SOTA WebPRM. In reward-guided trajectory search on WebArena-Lite, WebArbiter delivers gains of up to 6.4 points in success rate over the best prior WebPRM.
@@ 3 Methodology
In this section, we present the design of WebArbiter. Fig. 2 provides an overview of the WebArbiter framework, including the two-stage training pipeline and the inference-time principle-guided decision process. We begin by framing web navigation as a Partially Observable Markov Decision Process (POMDP) in §3.1, then describe how we construct a pairwise-preference dataset for training in §3.2. We introduce the training pipeline of WebArbiter model in §$3.$ For clarity, we summarize all notations in Appendix A.
@@ 3.1 Background
We formalize web navigation as a POMDP. The environment $\mathcal{E}$ is defined by a state space $\mathcal{S}$, an action space $\mathcal{A}$, and an observation space $\mathcal{O}$. $T:\mathcal{S}\!\times\!\mathcal{A}\!\to\!\mathcal{S}$ denotes the state transition function. At step $p$, the agent receives a partial observation $o_{p}\!\in\!\mathcal{O}$, executes $a_{p}\!\in\!\mathcal{A}$, and transitions to $s_{p+1}=T(s_{p},a_{p})$ with a new observation $o_{p+1}$. Following WebArena (Zhou et al., 2024), we represent observations using accessibility trees, i.e., text-only encodings of visible interactive elements and their attributes. Given a task instruction $\mathcal{I}$ and the initial state $s_{0}\in\mathcal{S}$, the agent aims to generate a trajectory $\tau=(a_{1},\dots,a_{P})$ that completes the task. Here $P$ is the trajectory length and $a_{p}\in\mathcal{A}$ denotes the action at step pp. The task evaluator determines whether the task is completed based on the final state.
@@ 4 WebPRMBench
Environment Mind2Web WebArena AssistantBench WorkArena Total
@@ 4.1 Benchmark Construction
For each state, the action from the successful trajectory is retained as the positive label, and four rejected alternatives with associated reasoning traces are synthesized to form preference pairs. To ensure data quality, we sample negatives from diverse policy models to broaden coverage, apply rule-based filters to remove invalid or mismatched actions, discard inconsistent cases, and conduct expert verification to ensure reliability. We also conduct targeted auditing to eliminate false negatives. To avoid positional bias, the positive action is not fixed to a specific side and may appear on either side of the preference pair. Reasoning traces are truncated to a fixed length to minimize formatting noise. The resulting benchmark spans 1,150 step-level preference instances across four environments, as shown in Tab. 1. Full construction details and benchmark statistics are provided in Appendix E.
@@ 5.1.2 Main Results
WebArbiter Significantly Outperforms Baselines. As shown in Tab. 2, WebArbiter achieves the highest Avg. Pairwise Acc and Avg. BoN Acc, surpassing both proprietary and open-source LLMs. While LLM-as-judge methods often maintain moderate Pairwise Acc, their performance drops sharply on BoN Acc, revealing poor robustness to hard negatives. In contrast, WebArbiter sustains strong results on both metrics, establishing its reliability under realistic multi-candidate settings. We show that the proposed training pipeline generalizes across backbone families in Appendix G, where WebArbiter-8BQwen3 achieves the highest Avg. BoN Acc of 76.66%, and provide additional analyses of inference-time scaling in Appendix H.
Advantage over the SOTA WebPRM. WebShepherd (Chae et al., 2025) represents the previous SOTA WebPRMs. Trained on the same WEBPRM Collection, which was drawn from the Mind2Web environment, WebArbiter-7B achieves an Avg. BoN Acc of 74.60%, surpassing WebShepherd-8B by an absolute gain of 31%. Unlike WebShepherd, which relies on fragile checklists, WebArbiter employs principle-guided reasoning, yielding judgments robust to environment and page variations. Case studies illustrating these differences are provided in Appendix I.
@@ 5.2.1 Experimental Setup and Implementations
Reward-guided trajectory search represents one of the most practical applications of PRMs, as it directly leverages fine-grained step-level supervision to improve decision quality during agent execution. To evaluate WebArbiter in this setting, we conduct experiments on WebArena-Lite (Liu et al., 2024b), which contains diverse, long-horizon tasks such as online shopping and content management, closely reflecting real-world web activities. Performance is measured with Success Rate. Following WebShepherd (Chae et al., 2025), we adopt a Best-of-N sampling strategy: the policy model generates $N=5$ candidate actions for each step, and WebArbiter selects the most promising one through a Knockout Tournament mechanism (Guo et al., 2025b). We evaluate two policies, GPT-4o-mini (OpenAI, 2024a) and GPT-4o (OpenAI, 2024b).
@@ 5.2.2 Downstream Analysis across Domains
As shown in Tab. 5, WebArbiter achieves substantial average improvements under both policy models, significantly outperforming all baselines. These gains stem from two main factors. First, reasoning mitigates spurious correlations that often mislead WebPRMs in domains such as Shopping and Reddit. The improvements on Shopping are particularly pronounced, as these tasks require dense semantic retrieval and inference: stronger policies can propose more promising candidate actions, and WebArbiter’s structured reward modeling further amplifies these advantages. Second, in GitLab, tasks frequently admit multiple equivalent paths. WebShepherd is brittle under such variability, whereas WebArbiter reasons over historical trajectories and the current state to evaluate action validity, enabling stronger generalization in dynamic workflows. We provide qualitative case studies in Appendix I to further illustrate these failure modes of checklist-based supervision. MAP presents a distinct pattern: tasks primarily involve information retrieval over OpenStreetMap, where success hinges on the semantic precision of search queries. Naïve reward signals fail to assess whether a query targets the intended geographic entity, while WebArbiter nearly doubles the baseline by reasoning over query-task alignment. By contrast, CMS exhibits a more template-driven structure, where actions closely follow standardized patterns. In such settings, checklist-based supervision remains comparatively effective, which narrows the relative performance gap. Overall, WebArbiter’s reasoning-first design consistently provides robust, interpretable, and scalable supervision across diverse domains.
@@
Policy WebPRM Shopping CMS Reddit GitLab MAP Avg. $\Delta$
GPT-4o-mini w/o Trajectory Search* 21.74 22.86 19.05 34.38 19.35 23.48
GPT-4o-mini 24.44 22.86 26.32 33.33 15.38 24.47 +0.99
WebShepherd-8B* 26.09 45.71 23.81 40.62 35.48 34.34 +10.86
$\bigstar$ WebArbiter-7B 37.78 42.86 36.84 46.67 38.46 40.52 +17.04
GPT-4o w/o Trajectory Search* 23.91 31.43 28.57 56.25 19.35 31.90
GPT-4o-mini 26.67 37.14 42.11 40.00 19.23 33.03 +1.13
WebShepherd-8B* 30.43 42.86 47.62 46.88 35.48 40.65 +8.75
$\bigstar$ WebArbiter-7B 44.44 42.86 52.63 56.67 38.46 47.01 +15.11
@@ 6 Conclusion
We presented WebArbiter, a reasoning-first, principle-inducing process reward model that frames reward modeling as structured text generation and produces auditable step-level judgments with rationales. Through reasoning distillation and RL, WebArbiter converts superficial correlations into robust, progress-aware signals that verify task advancement, yield consistent step-level judgments across trajectories, and generalize across dynamic web environments. To support systematic evaluation, we released WebPRMBench, the first comprehensive evaluation benchmark spanning diverse environments for WebPRMs in web navigation, covering four domains with diverse tasks and step-level preference annotations. Experiments demonstrate SOTA performance on WebPRMBench and improvements in reward-guided trajectory search on WebArena-Lite, establishing principle-guided reasoning WebPRMs as a robust and interpretable foundation for scalable web agents.
Limitations and Future Work
WebArbiter relies on text-based accessibility tree representations without visual observations; in environments where layout, spatial arrangement, or visual cues carry task-relevant information, this text-only formulation may miss critical signals for accurate progress assessment. We provide concrete failure cases illustrating this limitation in Appendix J. Incorporating multimodal observations into the reward modeling process is a natural extension. Training and evaluation are also conducted exclusively in English-language web environments, where multilingual settings may introduce additional challenges such as divergent page structures and interaction conventions. Finally, our current use of WebPRMs focuses on inference-time trajectory search; investigating whether PRM-based step-level supervision can serve as a dense reward signal during agent post-training, thereby improving the policy itself, is an important open direction.
@@ Appendix A Notation Summary
  • $s_{p},o_{p},a_{p}$: state, observation, and action at step pp.
@@
  • $\mathcal{D}_{\text{Bench}}$: evaluation dataset for WebPRMBench.
@@
  • $Q$: number of distractors in WebPRMBench (set to 44).
@@ Appendix C Training Details
All training is conducted on 8 NVIDIA A100-80GB GPUs with fixed random seeds. Our training framework is based on LLama-Factory (Zheng et al., 2024) and VERL (Sheng et al., 2024). In addition to the Qwen2.5 (Qwen et al., 2025) backbones used in the main experiments, we also train WebArbiter on Qwen3 (Yang et al., 2025) backbones (4B and 8B); results are reported in Appendix G.
Distillation Stage. For Qwen2.5 backbones (3B and 7B), we use a learning rate of 8e-4 with LoRA rank 128. For Qwen3-4B, the learning rate is 1e-3 with LoRA rank 96. For Qwen3-8B, the learning rate is 1e-3 with LoRA rank 64 and a reduced batch size of 64. All variants use a cosine learning rate scheduler with a warmup ratio of 0.1 and a maximum sequence length of 8,192 tokens. Models are trained for 5 epochs.
RLVR Stage. We employ the VERL framework for GRPO training. For Qwen2.5 backbones, the learning rate is set to 7e-6 (7B) and 9e-6 (3B). For Qwen3 backbones (4B and 8B), the learning rate is 1e-5. The training uses a fixed batch size of 512 with a mini-batch size of 128, and adopts Fully Sharded Data Parallel (FSDP) for enhanced memory efficiency. For rollout generation, we deploy vLLM with tensor parallelism of 4 and GPU memory utilization limited to 0.4. Response sampling uses standard parameters (temperature=1.0, top-p=1.0). The rollout group size is $G=7$ for Qwen2.5 backbones and $G=12$ for Qwen3 backbones. We apply KL regularization with a coefficient of $1.0\times 10^{-3}$ and a clip ratio of 0.2. The maximum input sequence length is 8,192 tokens, and the maximum response length is 4,096 tokens.
@@ E.2 Dataset Composition and Statistics
Appendix G Generalization Across Backbone Families
To verify that the proposed two-stage training pipeline is not tied to a specific model family, we train WebArbiter on Qwen3 (Yang et al., 2025) backbones (4B and 8B) in addition to the Qwen2.5 backbones reported in the main paper. All variants use the same training data, distillation strategy, and RL procedure; only backbone-specific hyperparameters differ (see Appendix C).
Table 8: Results on WebPRMBench with additional Qwen3 backbones. Models marked with $\bigstar$ are ours.
Models Mind2Web WebArena AssistantBench WorkArena Avg.
Pairwise BoN Pairwise BoN Pairwise BoN Pairwise BoN Pairwise BoN
3~4B
$\bigstar$ WebArbiter-3BQwen2.5 93.32 78.42 81.97 56.22 78.33 46.67 81.01 54.81 83.65 59.06
$\bigstar$ WebArbiter-4BQwen3 98.55 94.73 83.21 61.19 92.50 83.33 76.68 50.96 87.73 72.55
7~8B
$\bigstar$ WebArbiter-7BQwen2.5 97.07 89.53 88.43 68.66 89.17 70.00 82.09 70.19 89.19 74.60
$\bigstar$ WebArbiter-8BQwen3 98.33 94.09 86.92 67.16 92.50 80.00 86.66 65.38 91.10 76.66
As shown in Tab. 8, WebArbiter achieves strong performance on both backbone families. Within the 3~4B group, WebArbiter-4BQwen3 substantially outperforms WebArbiter-3BQwen2.5 across all environments, improving Avg. BoN Acc from 59.06% to 72.55%. This result approaches WebArbiter-7BQwen2.5 (74.60%) with roughly half the parameters, suggesting that stronger base models amplify the benefits of principle-guided reasoning distillation. Within the 7~8B group, WebArbiter-8BQwen3 achieves the highest Avg. BoN Acc of 76.66%, outperforming WebArbiter-7BQwen2.5 by 2.06 points.
Overall, these results confirm that the proposed two-stage training pipeline generalizes across model families and benefits from stronger base models without requiring pipeline modifications.
Appendix H Inference-Time Scaling